# Sparsity-Promoting Reachability Analysis and Optimization of Constrained Zonotopes

### v2

@@ Abstract
The constrained zonotope is a polytopic set representation widely used for set-based analysis and control of dynamic systems. This paper develops methods to formulate and solve optimization problems for dynamic systems in real time using constrained zonotope reachability analysis. An alternating direction method of multipliers (ADMM) algorithm is presented that makes efficient use of the constrained zonotope structure. To increase the efficiency of the ADMM iterations, reachability calculations are presented that increase the sparsity of the matrices used to define a constrained zonotope when compared to typical methods. The developed methods are used to formulate and solve predictive control, state estimation, and safety verification problems. Numerical results show that optimization times using the proposed approach are competitive with state-of-the-art QP solvers and conventional problem formulations. A combined set-valued state estimation and moving horizon estimation algorithm is presented and experimentally demonstrated in the context of robot localization.
@@ I Introduction
Novel set-theoretic methods continue to be developed as important tools for analyzing and controlling dynamic systems. For example, sets can be used to rigorously bound the reachable states of a system or to define constraints to which a controller must adhere. Zonotopes and generalizations of zonotopes have become particularly popular set representations, as is evidenced by their appearance in several open-source toolboxes [1, 2, 3, 4, 5, 6, 7, 8]. Key to the appeal of these set representations is that they have closed-form expressions for many important set operations and, in the context of reachability analysis, have favorable complexity growth when compared to traditional set representations such as halfspace representation (H-rep) polytopes, vertex representation (V-rep) polytopes, and ellipsoids [9].
These attributes make zonotopes and generalizations of zonotopes attractive candidates for use in real-time reachability analysis. In [10], for example, zonotope-based reachability analysis was used to verify the safety of system trajectories in autonomous driving experiments. In addition to safety verification, reachability analysis can be used to formulate optimal control and planning problems, as in [11, 12, 13, 14].
@@
I-A Gaps in the Literature
Despite the favorable complexity growth of constrained zonotopes when compared to more traditional set representations, sets with high complexity, e.g., those defined by hundreds or thousands of constraints and generator variables, are often produced in practice [18, 22, 17]. High complexity sets may have large memory requirements and generally take longer to analyze using numerical optimization algorithms than lower complexity sets. This challenge has motivated work on complexity reduction techniques [19, 15, 23]. Complexity is reduced by finding sets that are equivalent to or approximate the original set, but for which the matrices and vectors used to define the set have lower dimension. Sparse matrices can also be used to reduce the memory demands of a constrained zonotope representation, as pointed out in [19]. The sparsity patterns of these matrices have not been explicitly considered, however.
In the specific contexts of motion planning and moving horizon estimation, the particular properties and structure of hybrid zonotopes are exploited within a mixed-integer optimization routine in [30, 31, 32]. While some of the optimization techniques presented in these papers apply to constrained zonotopes, the methods are restricted to multi-stage problem structures and problems with diagonal cost matrices. Efficient solution of more general optimization problems formulated using constrained zonotopes has not been considered.
I-B Contributions
This paper considers the real-time formulation and solution of constrained optimization problems for dynamic systems. Our approach is illustrated in Fig. 1, where the feasible space of a constrained optimal control or estimation problem for a discrete-time dynamic system is constructed using constrained zonotope reachability analysis, and then ADMM is used to find the optimal trajectory within the reachable sets. Expressing the feasible space of an optimization problem through set-theoretic operations provides a unique framework from which to consider its structure and computational complexity. This leads to new insights in constructing a problem representation that is amenable to efficient optimization, and in selecting an optimization algorithm to most efficiently exploit the problem representation. We pursue computational efficiencies in both the reachability and optimization steps to facilitate their real-time execution.
The key contributions of this paper are as follows: (1) Novel reachability calculations for affine dynamics using constrained zonotopes decrease the number of non-zero elements in the matrices that define the set when compared to existing methods. (2) An ADMM formulation that makes efficient use of the constrained zonotope structure solves convex quadratic programs (QPs) for which the constraint set is represented as a constrained zonotope. (3) A method is proposed for efficiently verifying that a constrained zonotope is empty using the ADMM iterations. (4) An algorithm based on reachability analysis of constrained zonotopes combines moving horizon estimation (MHE) and SVSE. The effectiveness of this algorithm is demonstrated experimentally.
The new methods proposed in this paper are available in an open-source toolbox that facilitates reachability calculations using constrained zonotopes and sparse linear algebra with tailored optimization routines. This toolbox is implemented in C++ with Python bindings and is available at https://github.com/psu-PAC-Lab/ZonoOpt.
@@ I-C Outline
The remainder of this paper is organized as follows. Preliminary information is given in Sec. II. In Sec. III, reachability calculations for linear systems using constrained zonotopes are presented that result in matrices with high sparsity. Sec. IV proposes an ADMM algorithm that is tailored to constrained zonotopes. In Sec. V, the sparse reachability calculations and ADMM algorithm are applied to MPC, MHE, and safety verification problems. Numerical results for these problems are given in Sec. VI. Using the proposed approach based on sparse constrained zonotope reachability analysis and ADMM, the MPC problems are solved faster than when using state-of-the-art QP solvers and a commonly used sparse MPC problem formulation, especially for large problem sizes. The SVSE/MHE example illustrates an equivalence between reachability analysis and MHE, as a set-valued state estimate is extracted directly from the MHE problem formulation. The infeasibility certificate is used for efficient safety verification of system trajectories in a third numerical example. Sec. VII experimentally demonstrates the real-time capabilities of the proposed methods, with the combined SVSE/MHE algorithm applied for online robot localization. Sec. VIII concludes the paper.
@@ II-B Zonotopes and Constrained Zonotopes
where (3a) is the affine map, (3b) is the Cartesian product, (3c) is the Minkowski sum, and (3d) is the generalized intersection.
@@ II-C Alternating Direction Method of Multipliers
$\text{s.t.}\;A\mathbf{x}+B\mathbf{z}=\mathbf{c}\;.$ (4b)
@@
Subject to mild conditions on (4), the iterates (5) are guaranteed to converge to the optimal solution for feasible optimization problems. Typical stopping criteria are to monitor the primal and dual residuals $\mathbf{r}_{p}\in^{p}$ and $\mathbf{r}_{d}\in^{n}$ for convergence, where
@@ III Reachability with Constrained Zonotopes via Sparsity-Promoting Operations
For optimization algorithms that rely on sparse linear algebra, the computational complexity of optimizing over or assessing the feasibility of a constrained zonotope will depend on the number of non-zero elements of the matrices involved [35]. Furthermore, assessments of the computational complexity of set operations on constrained zonotopes have assumed dense linear algebra [9]. The following presents closed-form set operations on constrained zonotopes to compute reachable sets that result in matrices with significantly fewer non-zero elements than existing methods.
@@
where $\mathcal{S}$ and $\mathcal{U}$ are the state and input domain sets of the system dynamics, respectively [36]. The state domain set $\mathcal{S}$ captures limits on the system states and is particularly useful when using reachability analysis to build MPC and MHE problems, as demonstrated in Secs. V-A and V-B.
@@ III-A1 Standard Method
$\mathcal{X}_{k+1}=\left(A\mathcal{X}_{k}\oplus B\mathcal{U}\right)\cap S\;,\forall k\in\{0,...,N-1\}\;.$ (10)
@@
Constrained zonotopes are a popular choice for computing $\mathcal{X}_{N}$ because all of the operations in (10) have closed-form expressions as given in (3). Reachability analysis using constrained zonotopes is known to have favorable complexity growth when compared to methods based on other polytopic set representations [9]. However, these analyses do not consider the sparsity structure of the resulting sets. Indeed the sparsity of the $G$ and $A$ matrices in the $N$-step reachable set $\mathcal{X}_{N}$ can be improved significantly, as will be demonstrated in Secs. III-A2 and III-A3.
@@
From inspection of (11), the sparsity of matrices $G$ and $A$ in $\mathcal{X}_{N}=\left\langle G,\cdot,A,\cdot\right\rangle$ scales as $\operatorname*{\mathrm{nnz}}{(G)}\propto N$ and $\operatorname*{\mathrm{nnz}}{(A)}\propto N^{2}$. Define the number of generators and constraints for sets $\mathcal{X}_{0}$, $\mathcal{S}$, and $\mathcal{U}$ to be $(n_{G,0},n_{C,0})$, $(n_{G,s},n_{C,s})$, and $(n_{G,u},n_{C,u})$, respectively. The numbers of generators and constraints for $\mathcal{X}_{N}$ are then $n_{G}=n_{G,0}+N(n_{G,s}+n_{G,u})$ and $n_{C}=n_{C,0}+N(n_{C,s}+n_{C,u}+n_{x})$ where $n_{x}$ is the number of states for the system.
@@ III-A2 Graph of Function Method
$\Psi=\left\{\begin{bmatrix}\mathbf{p}&\mathbf{q}\end{bmatrix}^{T}\middle|\mathbf{p}\in\mathcal{D},\mathbf{q}\in\psi(\mathbf{q})\subseteq\mathcal{Q}\right\}\;,$ (12)
@@
Defining the rows of the generator matrix in (15) to be $G_{\psi,x,k}$, $G_{\psi,u,k}$, and $G_{\psi,x,k+1}$ and the columns of the constraint matrix to be $A_{\psi,x,k}$, $A_{\psi,u,k}$, and $A_{\psi,x,k+1}$, the generator and constraint matrices for the iterates (14) are given explicitly as
$G_{x,k+1}=\begin{bmatrix}G_{\psi,x,k+1}&0&0\end{bmatrix}\;$, (16a)
$A_{x,k+1}=\begin{bmatrix}A_{\psi,x,k}&A_{\psi,u,k}&A_{\psi,x,k+1}&0&0\\ 0&0&0&A_{x,k}&0\\ 0&0&0&0&A_{u}\\ G_{\psi,x,k}&0&0&-G_{x,k}&0\\ 0&G_{\psi,u,k}&0&0&-G_{u,k}\end{bmatrix}\;$. (16b)
In contrast with (11), the $N$-step reachable set $\mathcal{X}_{N}$ computed using (16) has constant $\operatorname*{\mathrm{nnz}}{(G)}$ with respect to $N$ and $\operatorname*{\mathrm{nnz}}{(A)}\propto N$. The numbers of generators and constraints for $\mathcal{X}_{N}$ using this method are $n_{G}=n_{G,0}+2N(n_{G,s}+n_{G,u})$ and $n_{C}=n_{C,0}+N(2n_{C,s}+n_{C,u}+2n_{x}+n_{u})$, where $n_{u}$ is the number of inputs to the system.
@@ III-A3 Sparsity-Promoting Method
For the $N$-step reachable set $\mathcal{X}_{N}$, the number of non-zero elements in $G$ is constant, i.e., $\operatorname*{\mathrm{nnz}}{(G)}=\operatorname*{\mathrm{nnz}}{(G_{s})}\;\forall N$, and $\operatorname*{\mathrm{nnz}}{(A)}\propto N$.
Equations (11) and (18) can be seen to be equivalent from the equality constraints. Further, the dimensions of the generator and equality constraint matrices $G$ and $A$ are the same in both cases. The matrices in (18), however, have much higher sparsity. When compared to (14), (17) has fewer generators and constraints. In both cases, $G$ is constant with respect to $N$ and $A\propto N$, but (18) has fewer non-zero elements than (16) by inspection.

### v3

# arXiv:2504.03885v3[eess.SY] 01 Mar 2026

@@ Abstract
The constrained zonotope is a polytopic set representation widely used for set-based analysis and control of dynamic systems. This paper develops methods to formulate and solve optimization problems for dynamic systems in real time using reachability analysis with constrained zonotopes. An alternating direction method of multipliers (ADMM) algorithm is presented that makes efficient use of the constrained zonotope structure. To increase the efficiency of the ADMM iterations, reachability calculations are presented that increase the sparsity of the matrices used to define a constrained zonotope when compared to typical methods. The developed methods are used to formulate and solve predictive control, state estimation, and safety verification problems. Numerical results show that optimization times using the proposed approach are competitive with state-of-the-art QP solvers and conventional problem formulations. A combined set-valued state estimation and moving horizon estimation algorithm is presented and experimentally demonstrated in the context of robot localization.
@@ I Introduction
Novel set-theoretic methods continue to be developed as important tools for analyzing and controlling dynamic systems. For example, sets can be used to rigorously bound the reachable states of a system or to define constraints to which a controller must adhere. Traditional set representations such as halfspace representation (H-rep) polytopes, vertex representation (V-rep) polytopes, ellipsoids, and intervals remain widely used. More modern offerings include include AH-polytopes [42], ellipsotopes [23], constrained convex generators [48], and zonotopic sets (i.e., zonotopes and generalizations of zonotopes). Zonotopic sets have become particularly popular, as evidenced by their appearance in several open-source toolboxes [21, 4, 55, 7, 25, 17, 43, 35]. Key to the appeal of these set representations is that they have closed-form expressions for many important set operations and have favorable complexity growth when compared to traditional set representations [3].
These attributes make zonotopic sets attractive candidates for use in real-time reachability analysis. In [2], for example, zonotope-based reachability analysis was used to verify the safety of system trajectories in autonomous driving experiments. In addition to safety verification, reachability analysis can be used to formulate optimal control and planning problems [46, 24, 11, 44].
@@
I-A Gaps in the Literature
Despite the favorable complexity growth of constrained zonotopes when compared to more traditional set representations, sets with high complexity, e.g., those defined by hundreds or thousands of constraints and generator variables, are often produced in practice [54, 58, 34]. High complexity sets may have large memory requirements and generally take longer to analyze using numerical optimization algorithms than lower complexity sets. This challenge has motivated work on complexity reduction techniques [31, 45, 36]. Complexity is reduced by finding sets that are equivalent to or approximate the original set, but for which the matrices and vectors used to define the set have lower dimension. Sparse matrices can also be used to reduce the memory demands of a constrained zonotope representation, as noted in [31]. The sparsity patterns of these matrices have not been explicitly considered, however.
In the specific contexts of motion planning and moving horizon estimation, the particular properties and structure of hybrid zonotopes—an extension of constrained zonotopes that includes integer variables—are exploited within a mixed-integer optimization routine in [39, 40, 53]. While some of the optimization techniques presented in these papers apply to constrained zonotopes, the methods are restricted to multi-stage problem structures and problems with diagonal cost matrices.
I-B Contributions
This paper considers the real-time formulation and solution of constrained optimization problems for dynamic systems. Our approach is illustrated in Fig. 1. The feasible space of a constrained optimal control or estimation problem for a discrete-time dynamic system is constructed using reachability analysis with constrained zonotopes, and then ADMM is used to find the optimal trajectory within the reachable sets. Expressing the feasible space of an optimization problem through set-theoretic operations provides a unique framework from which to consider its structure and computational complexity. This leads to new insights in constructing a problem representation that is amenable to efficient optimization, and in selecting an optimization algorithm to most efficiently exploit the problem representation. We pursue computational efficiencies in both the reachability and optimization steps to facilitate their real-time execution.
The key contributions of this paper are as follows:
  1. 1.
    Novel reachability calculations for affine dynamics using constrained zonotopes decrease the number of non-zero elements in the matrices that define the set when compared to existing methods.
  2. 2.
    An ADMM formulation makes efficient use of the structure of constraint sets represented as constrained zonotopes when solving convex quadratic programs (QPs).
  3. 3.
    A method is proposed for efficiently verifying that a constrained zonotope is empty using the ADMM iterations.
  4. 4.
    An algorithm based on reachability analysis of constrained zonotopes combines moving horizon estimation (MHE) and SVSE. The effectiveness of this algorithm is demonstrated experimentally.
The new methods proposed in this paper are available in an open-source toolbox [^1] https://github.com/psu-PAC-Lab/ZonoOpt, implemented in C++ with Python bindings, that facilitates reachability calculations using constrained zonotopes and sparse linear algebra with tailored optimization routines.
@@ I-C Outline
The remainder of this paper is organized as follows. Preliminary information is given in Sec. II. In Sec. III, reachability calculations for linear systems using constrained zonotopes are presented that result in matrices with high sparsity. Sec. IV proposes an ADMM algorithm that is tailored to constrained zonotopes. In Sec. V, the sparse reachability calculations and ADMM algorithm are applied to MPC, MHE, and safety verification problems. Numerical results for these problems are given in Sec. VI. Using the proposed approach based on reachability analysis with sparse constrained zonotopes and ADMM, the MPC problems are solved faster than when using state-of-the-art QP solvers and a commonly used sparse MPC problem formulation, especially for large problem sizes. The SVSE/MHE example illustrates an equivalence between reachability analysis and MHE, as a set-valued state estimate is extracted directly from the MHE problem formulation. The infeasibility certificate is used for efficient safety verification of system trajectories in a third numerical example. Sec. VII experimentally demonstrates the real-time capabilities of the proposed methods, with the combined SVSE/MHE algorithm applied for online robot localization. Sec. VIII concludes the paper.
@@ II-B Zonotopes and Constrained Zonotopes
where (3a) is the affine map $\{R\mathbf{z}+\mathbf{s}|\mathbf{z}\in\mathcal{Z}\}$, (3b) is the Cartesian product $\{[\mathbf{z}_{1}^{T}~\mathbf{z}_{2}^{T}]^{T}|\mathbf{z}_{1}\in\mathcal{Z}_{1},\;\mathbf{z}_{2}\in\mathcal{Z}_{2}\}$, (3c) is the Minkowski sum $\{\mathbf{z}_{1}+\mathbf{z}_{2}|\mathbf{z}_{1}\in\mathcal{Z}_{1},\;\mathbf{z}_{2}\in\mathcal{Z}_{2}\}$, and (3d) is the generalized intersection, $\{\mathbf{z}\in\mathcal{Z}_{1}|R\mathbf{z}\in\mathcal{Z}_{2}\}$.
@@ II-C Alternating Direction Method of Multipliers
$\text{s.t.}\;A\mathbf{x}+B\mathbf{z}=\mathbf{c}\;,$ (4b)
where $\mathbf{x}\in^{n}$, $\mathbf{z}\in^{m}$, and $\mathbf{c}\in^{p}$, and $f:^{n}\rightarrow\cup\{\infty\}$ and $g:^{m}\rightarrow\cup\{\infty\}$
@@
The convergence of (5) to feasible and optimal $\mathbf{x}^{*}$, $\mathbf{z}^{*}$, and $\mathbf{u}^{*}$ is guaranteed when 1) $f$ and $g$ are closed, proper, and convex, and 2) the augmented Lagrangian
[…]
has a saddle point [10].
Typical stopping criteria are to monitor the primal and dual residuals $\mathbf{r}_{p}\in^{p}$ and $\mathbf{r}_{d}\in^{n}$ for convergence, where
@@ III Reachability with Constrained Zonotopes via Sparsity-Promoting Operations
For optimization algorithms that rely on sparse linear algebra, the computational complexity of optimizing over or assessing the feasibility of a constrained zonotope will depend on the number of non-zero elements of the matrices involved [13]. However, previous assessments of the computational complexity of set operations on constrained zonotopes have assumed dense linear algebra [3]. The following presents closed-form set operations on constrained zonotopes for estimation and control applications that result in matrices with significantly fewer non-zero elements than existing methods.
@@
where $\mathcal{S}$ and $\mathcal{U}$ are the state and input domain sets of the system dynamics, respectively [47]. The state domain set $\mathcal{S}$ captures limits on the system states. This is particularly useful in SVSE and when using reachability analysis to build MPC and MHE problems. In applications where there is no a priori bound $\mathcal{S}$ on the system states, intersections are not necessary and reachability analysis with zonotopes can be used.
@@ III-A Methods for Computing the N-Step Reachable Set
We next present three different methods for computing the N-step reachable set of the system in (10), each of which achieves a different sparsity in the matrices defining the reachable set.
@@ III-A1 Standard Method
$\mathcal{X}_{k+1}=\left(A\mathcal{X}_{k}\oplus B\mathcal{U}\right)\cap{\mathcal{S}}\;,\forall k\in\{0,...,N-1\}\;.$ (11)
@@ III-A2 Graph of Function Method
$\Psi=\left\{\begin{bmatrix}\mathbf{p}&\mathbf{q}\end{bmatrix}^{T}\middle|\mathbf{p}\in\mathcal{D},\mathbf{q}{=}\psi({\mathbf{p}}),{\mathbf{q}}\subseteq\mathcal{Q}\right\}\;,$ (13)
@@
Defining the rows of the generator matrix in (16) to be $G_{\psi,1}$, $G_{\psi,2}$, and $G_{\psi,3}$ and the columns of the constraint matrix to be $A_{\psi,1}$, $A_{\psi,2}$, and $A_{\psi,3}$, the generator and constraint matrices for the iterates (15) are given explicitly as
$G_{x,k+1}=\begin{bmatrix}{G_{\psi,3}}&0&0\end{bmatrix}\;$, (17a)
$A_{x,k+1}=\begin{bmatrix}{A_{\psi,1}}&{A_{\psi,2}}&{A_{\psi,3}}&0&0\\ 0&0&0&A_{x,k}&0\\ 0&0&0&0&A_{u}\\ {G_{\psi,1}}&0&0&-G_{x,k}&0\\ 0&{G_{\psi,2}}&0&0&-G_{u}\end{bmatrix}\;$. (17b)
@@ III-A3 Sparsity-Promoting Method
Equations (12) and (19) can be seen to be equivalent from the constrained zonotope equality constraints.
TABLE I: Memory complexity of the $N$-step reachable set $\mathcal{X}_{N}$ using the standard (11), graph of function (15), and sparsity-promoting (18) methods.
[…]
Eq. (11) $N(n_{G,s}+n_{G,u})$ $N(n_{C,s}+n_{C,u}+n_{x})$ $N(n_{x}n_{G,u})$ $N^{2}(\frac{n_{x}n_{G,u}}{2})+N((\frac{n_{x}}{2}+n_{C,u})n_{G,u}$
$n_{G,0} +$$n_{C,0} +$$n_{x}n_{G,0} +$$(n_{x}+n_{C,s})n_{G,s}+n_{x}n_{G,0})+n_{C,0}n_{G,0}$
Eq. (15) $2N(n_{G,s}+n_{G,u})$ $N(2n_{C,s}+n_{C,u}+2n_{x}+n_{u})$ $n_{x}(n_{G,s}+n_{G,u})$ $N((3n_{x}+2n_{C,s})n_{G,s}$
$n_{G,0} +$$n_{C,0} +$$(2n_{u}+2n_{C,u}+2n)n_{G,u})+n_{C,0}n_{G,0}$
Eq. (18) $N(n_{G,s}+n_{G,u})$ $N(n_{C,s}+n_{C,u}+n_{x})$ $n_{x}n_{G,s}$ $N((2n_{x}+n_{C,s})n_{G,s}+(n_{x}+n_{C,u})n_{G,u})$
$n_{G,0} +$$n_{C,0} +$$n_{C,0}n_{G,0}$
III-A4 Memory Complexity of the N-Step Reachable Set
Table I gives the memory complexity for computing the set $\mathcal{X}_{N}$ with the standard, graph of function, and sparsity-promoting methods. The number of generators and constraints for sets $\mathcal{X}_{0}$, $\mathcal{S}$, and $\mathcal{U}$ are $(n_{G,0},n_{C,0})$, $(n_{G,s},n_{C,s})$, and $(n_{G,u},n_{C,u})$, respectively. The number of states and inputs are $n_{x}$ and $n_{u}$. Upper bounds for the number of non-zero elements in a matrix are denoted as $\overline{\operatorname*{\mathrm{nnz}}}(\cdot)$. Bounds are computed by assuming that matrix products result in fully dense matrices. For clarity of exposition, we also assume that $G_{x,0}$, $A_{x,0}$, $G_{u}$, $A_{u}$, $G_{s}$, and $A_{s}$ are dense.
The standard and sparsity-promoting methods have the same number of generators and constraints, but the matrices $G_{N}$ and $A_{N}$ have much higher sparsity for the sparsity-promoting method. In particular, $\operatorname*{\mathrm{nnz}}{(G_{N})}\propto 1$ and $\operatorname*{\mathrm{nnz}}{(A_{N})}\propto N$ for (18), while $\operatorname*{\mathrm{nnz}}{(G_{N})}\propto N$ and $\operatorname*{\mathrm{nnz}}{(A_{N})}\propto N^{2}$ for (11). To see that $\operatorname*{\mathrm{nnz}}{(A_{N})}\propto N^{2}$ for (11), observe from (12b) that $\operatorname*{\mathrm{nnz}}{(A_{N})}=\operatorname*{\mathrm{nnz}}{(A_{N-1})}+\operatorname*{\mathrm{nnz}}{(AG_{N-1})}+\cdots$, and therefore $\operatorname*{\mathrm{nnz}}{(A_{N})}=\operatorname*{\mathrm{nnz}}{(A_{0})}+\sum_{i=1}^{N}\operatorname*{\mathrm{nnz}}{(AG_{i-1})}+\cdots\propto\sum_{i=1}^{N}(i-1)\propto\frac{N(N-1)}{2}$. When compared to the graph of function method, the sparsity-promoting method has fewer generators and constraints. The sparsity scaling is $\operatorname*{\mathrm{nnz}}{(G_{N})}\propto 1$ and $\operatorname*{\mathrm{nnz}}{(A_{N})}\propto N$ in both cases, but inspecting Table I shows that $\operatorname*{\mathrm{nnz}}{(G_{N})}$ and $\operatorname*{\mathrm{nnz}}{(A_{N})}$ are lower for (19) than (17).